# SAFR: Neuron Redistribution for Interpretability

### v1

@@ 6 Related Work
Elhage et al. (2022) illustrated superposition in simplified networks. Concurrently, theoretical explorations Scherlis et al. (2022); Henighan et al. (2023); Hänni et al. (2024); Gurnee et al. (2023); Marshall and Kirchner (2024); Hänni et al. (2024); Katta (2024) and applications in language models Chen et al. (2023, 2024) have expanded the field.
Interpretability efforts include works like Dreyer et al. (2024); Black et al. (2022); Wang et al. (2023). Challenges in lifelong knowledge editing Hu et al. (2024) and identifying universal feature spaces across models Lan et al. (2024) mark promising directions for future research.
@@ 7 Conclusion
In this work, we introduced SAFR, a novel approach to enhance model interpretability by strategically regularizing feature superposition. Experiments on SST-2 and IMDB show that SAFR improves interpretability, as measured by our SRS metric, without compromising prediction performance.
Our method provides insights into the relationship between superposition and interpretability in transformers and offers a framework for visualizing neuron allocation across layers. This work contributes to mechanistic interpretability and suggests promising directions for extending the approach to larger models and broader applications.
@@ 9 Ethic Statements
Appendix A Dataset Statistics
This section provides the statistical summary for the datsets and model.
@@ Appendix A Dataset Statistics
FFN1 (256, 1024)
FFN2 (1024, 256)
Table 5: Summary statistic for the model: Layer dimensions.
Appendix B Cross Entropy Loss
Given an input sequence of $n$ tokens $S=\langle s_{1},\ldots,s_{n}\rangle$ and a transformer $f(\cdot)$ with $L$ layers, let $h^{\ell}_{i}$ denote the hidden representation obtained at layer $\ell\in[1,\ldots,L]$ and position $i\in[1,\ldots,n]$. Following previous work Scherlis et al. (2022); Elhage et al. (2022), we define interference, polysemanticity, and capacity as follows. The baseline is defined via the original model using the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{m=1}^{M}\sum_{g=1}^{G}y_{m}^{g}\log(\hat{y}_{m}^{g})$
where $M$ is the number of samples in the dataset and $G$ is the number of classes in text classification.
Additionally, for the input sequence $S$, let:
  • $\hat{y}_{t}$ represent the predicted probability distribution over the vocabulary at position $t$,
  • $y_{t}$ represent the true one-hot encoded distribution for the token at position $t$,
  • $V$ denote the size of the vocabulary, and
  • $c_{t}$ be the index of the correct token at position $t$.
The cross-entropy loss at position $t$ is defined as:
$\mathcal{L}_{t}=-\sum_{i=1}^{V}y_{t,i}\log(\hat{y}_{t,i})$
Since $y_{t}$ is a one-hot encoded vector, this simplifies to:
$\mathcal{L}_{t}=-\log(\hat{y}_{t,c_{t}})$
The total cross-entropy loss for the sequence $S$ is the average over all tokens in the sequence:
$\mathcal{L}=\frac{1}{n}\sum_{t=1}^{n}\mathcal{L}_{t}=-\frac{1}{n}\sum_{t=1}^{n}\log(\hat{y}_{t,c_{t}})$
where $\hat{y}_{t,c_{t}}$ is the predicted probability of the correct token at position $t$.
@@ Appendix C Neuron Allocation Accross Layers
This section presents the observations regarding neuron allocation across the various layers.
Figure 4: Cross Layers Output-1: Capacity. Capacity for FFN1 is interpretable.
Figure 5: Cross Layers Output-2: Capacity. Capacity for FFN1 is interpretable.
Figure 6: Cross Layers Output: Interference. Interference for FFN1 is interpretable.
@@
Model $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS
Baseline 71.81 75.81 74.96 -3.15 80.88 74.31 76.92 3.96
$\lambda_{Imp}=0$, $\lambda_{Inter}=0$ 71.24 71.98 61.98 9.26 79.07 73.30 64.70 14.37
$\lambda_{Imp}=0$, $\lambda_{Inter}=0.01$ 71.98 72.04 57.29 14.69 78.38 74.12 61.55 16.83
$\lambda_{Imp}=0$, $\lambda_{Inter}=1$ 71.98 69.07 61.86 10.12 78.05 74.05 55.96 22.09
$\lambda_{Imp}=0$, $\lambda_{Inter}=100$ 62.66 78.56 81.25 -18.59 77.19 72.60 64.46 12.73
$\lambda_{Imp}=0.01$, $\lambda_{Inter}=0$ 73.47 69.70 59.35 14.12 78.63 74.48 50.42 28.21
$\lambda_{Imp}=0.01$, $\lambda_{Inter}=0.01$ 74.10 73.13 62.78 11.32 78.23 74.52 51.19 27.04
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=0.1$ 72.78 70.67 60.43 12.35 77.64 74.06 50.05 27.59
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=1$ 73.87 74.27 60.38 13.49 76.88 73.38 50.02 26.86
$\lambda_{Imp}=1$, $\lambda_{Inter}=0$ 71.47 72.16 55.35 16.12 74.93 71.57 50.00 24.93
$\lambda_{Imp}=1$, $\lambda_{Inter}=0.1$ 72.96 75.30 54.72 18.24 73.64 70.72 50.00 23.64
$\lambda_{Imp}=1$, $\lambda_{Inter}=1$ 71.58 73.70 56.89 14.69 74.10 70.89 50.00 24.10
$\lambda_{Imp}=10$, $\lambda_{Inter}=10$ 70.84 73.24 57.00 13.84 60.73 58.73 50.00 10.73
$\lambda_{Imp}=100$, $\lambda_{Inter}=0$ 59.86 64.95 51.57 8.29 52.04 51.81 50.09 1.95
$\lambda_{Imp}=100$, $\lambda_{Inter}=100$ 61.46 62.04 51.06 10.40 51.80 51.72 50.00 1.80
Table 6: Ablation Study on SST-2 and IMDB datasets. $\lambda$ choosing influence prediction accuracy and model interpretability.

### v2

@@ 6 Related Work
Elhage et al. (2022) illustrated superposition in simplified networks, while subsequent works explored its theoretical, empirical, and applications Scherlis et al. (2022); Henighan et al. (2023); Hänni et al. (2024); Gurnee et al. (2023); Marshall and Kirchner (2024); Hänni et al. (2024); Katta (2024); Chen et al. (2023, 2024).
Interpretability research includes works such as Dreyer et al. (2024); Black et al. (2022); Wang et al. (2023). Meanwhile, challenges in knowledge Hu et al. (2024) and identifying universal feature spaces across models Lan et al. (2024) mark promising directions for future research.
@@ 7 Conclusion
In this work, we introduced SAFR, an approach to enhance model interpretability by strategically regularizing feature superposition. Experiments on SST-2 and IMDB show that SAFR improves interpretability, as measured by our SRS metric, without compromising prediction performance.
Our method provides insights into the relationship between superposition and interpretability and offers a framework for visualizing neuron allocation. It contributes to mechanistic interpretability and suggests promising directions for extending the approach to larger models and wider applications.
@@ 9 Ethic Statements
Acknowledgements
We thank the anonymous reviewers for their valuable comments. We thank the Chili Lab at Rice for helpful discussions and suggestions throughout this research.
Appendix A VMASK Introduction
VMASK Chen and Ji (2020) is a variational word mask layer that is inserted into a neural text classifier and trained with the model. It learns to limit the flow of globally irrelevant or noisy word-level feature information to subsequent network layers, thus forcing the model to focus on the important features for prediction.
Appendix B Statistics
This section provides the statistical summaries of the datasets and the model.
@@ Appendix B Statistics
FC1 (256, 1024)
FC2 (1024, 256)
Table 5: Summary statistics for the model.
@@ Appendix C Neuron Allocation Accross Layers
This section presents the observations regarding neuron allocation across the various layers, as visualized in Figure 4 and 5.
Figure 4: Cross Layers Output: Capacity. VMASK layer uses the importance scores it detects, while the attention layer uses normalized attention scores. The original sentence is “Preposterous and tedious, Sonny is spiked with unintentional laughter that, unfortunately, occurs too infrequently to make the film even a guilty pleasure.”(negative)
Figure 5: Cross Layers Output: Interference. The attention layer uses the attention weight matrix. The original sentence is “Preposterous and tedious, Sonny is spiked with unintentional laughter that, unfortunately, occurs too infrequently to make the film even a guilty pleasure.”(negative)
@@
Model $\text{Acc}_{S}(\%)$ $\text{Acc}_{\tilde{S}^{(r)}}(\%)$ $\text{Acc}_{\tilde{S}^{(k)}}(\%)$ SRS $\text{Acc}_{S}(\%)$ $\text{Acc}_{\tilde{S}^{(r)}}(\%)$ $\text{Acc}_{\tilde{S}^{(k)}}(\%)$ SRS
Baseline 70.21 67.12 66.21 4.00 80.14 77.11 76.60 3.54
$\lambda_{Imp}=0$, $\lambda_{Inter}=0$ 72.56 69.47 64.44 8.12 78.43 76.10 73.56 4.87
$\lambda_{Imp}=0$, $\lambda_{Inter}=0.01$ 72.84 70.27 62.78 10.06 81.16 78.20 74.61 6.55
$\lambda_{Imp}=0$, $\lambda_{Inter}=1$ 73.01 69.07 60.32 12.69 80.44 76.89 71.84 8.60
$\lambda_{Imp}=0$, $\lambda_{Inter}=100$ 59.35 57.75 56.38 2.97 77.83 74.90 67.64 10.19
$\lambda_{Imp}=0.01$, $\lambda_{Inter}=0$ 72.67 69.87 61.41 11.26 80.00 77.02 70.98 9.02
$\lambda_{Imp}=0.01$, $\lambda_{Inter}=0.01$ 73.07 71.07 58.49 14.58 80.30 77.42 64.58 15.72
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=0.1$ 72.96 70.61 55.75 17.21 79.24 75.96 53.99 25.25
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=1$ 72.67 69.18 59.58 13.09 78.45 75.05 49.97 28.48
$\lambda_{Imp}=1$, $\lambda_{Inter}=0$ 71.30 66.96 55.06 16.24 74.95 72.30 50.07 24.88
$\lambda_{Imp}=1$, $\lambda_{Inter}=0.1$ 71.81 69.13 56.38 15.43 74.89 71.72 50.08 24.81
$\lambda_{Imp}=1$, $\lambda_{Inter}=1$ 71.93 70.67 56.78 15.15 74.78 71.44 53.94 20.84
$\lambda_{Imp}=100$, $\lambda_{Inter}=0$ 64.67 63.75 57.92 6.75 51.53 51.20 50.00 1.53
$\lambda_{Imp}=100$, $\lambda_{Inter}=100$ 63.01 61.64 54.20 8.81 53.48 53.33 50.00 3.48
Table 6: Ablation Study on SST-2 and IMDB datasets. The choice of $\lambda$ affects both prediction accuracy and model interpretability.